# Writing – review & editing, Writing – original draft, Visualization, Data curation ,Investigation, Formal analysis, Validation, Software, Methodology, Conceptualization

### v2

@@ 2.3 Indistinguishability in the Tabular Domain
Although there are many well-established anomaly detection techniques for the image domain, not all of them are directly applicable to tabular data. Since some methods fail to fully capture multivariate dependencies, selecting the right approach is crucial. Methods such as convolutional autoencoders (AEs) and generative adversarial networks (GANs) leverage spatial correlations but often struggle to capture the complex dependencies found in tabular data and therefore should be modified for the tabular domain [4]. Similarly, traditional statistical techniques like t-tests and the ANOVA assume normality and independence [16], limiting their ability to detect anomalies in the tabular domain where variables might not follow normal distributions and the relationships between variables are often complex and non-linear. Other simple approaches, such as Z-score thresholding, focus on individual feature deviations but fail to account for multivariate dependencies. Additionally, some commonly used anomaly detection techniques have inherent limitations for tabular data. For instance, PCA-based methods assume linear relationships, limiting their effectiveness in detecting non-linear dependencies in high-dimensional datasets [45].
@@
More suitable approaches for tabular data include AEs, which learn latent representations to detect deviations from expected patterns [45], and the isolation forest (IF) algorithm, which identifies anomalies by isolating points in a partitioned feature space. However, while effective for detecting extreme outliers, this algorithm is less sensitive to subtle structural inconsistencies [51].
@@
In our evaluation (see Section 4.5), we examined two decision-based attacks: the boundary attack [6] and the HopSkipJump attack [9], which were adjusted to address tabular constraints and evaluated by  [8]. We implemented these attacks by applying additional constraints based on feature distributions (e.g., limiting values according to statistical information) and making adjustments to ensure validity (e.g., [0,1] values for binary features and ordinal values for categorical features) and editability (guided by domain knowledge).
In addition, we performed transferability-based gradient attacks based on the architecture proposed in [36, 21], which is a sophisticated attack designed to craft valid and consistent adversarial samples.
@@
To ensure coherence and maintain feature consistency when crafting adversarial samples, we introduce a novel technique for perturbing dependent features. Specifically, for each dependent feature, we train a regression model on the benign samples, excluding the target feature. This model then predicts the value of the dependent feature based on the values of the other features in the sample.
During the adversarial attack process, when perturbations are applied to features, the regression model corresponding to each dependent feature is queried to predict its value, ensuring that the dependent feature remains consistent with the rest of the sample. This approach ensures that the perturbations respect indirect dependencies between features, thereby maintaining the overall coherence of the sample.
In our implementation of both query- and transferability-based attack strategies, the attack process includes an additional step in which the dependent features, as identified by domain knowledge, are adjusted according to the predictions from the regression models. This step helps maintain feature consistency and reduces the likelihood of generating detectable anomalies.
@@ 3.3 Evaluating the Quality of Adversarial Attacks
In the IF algorithm, a sample is considered anomalous if it is classified as an outlier. In the AE model, an anomaly is identified when its reconstruction error exceeds a threshold, defined in Eq. 1:
$\begin{split}Threshold(x_{c},\>AE_{c})=mean(test_{c,\>r.\,error})\,+\\ 0.5*std(test_{c,\>r.\,error})\end{split}$ (1)
where $x_{c}$ is a sample classified as class $c$, $AE_{c}$ is the autoencoder trained on benign samples of class $c$, and $test_{c,r.,error}$ represents the reconstruction error on a validation set of class $c$ during the AE’s training.
@@
The AE model consists of an embedding layer with 64 neurons, and it was trained using the Adam optimizer with a learning rate of $31e-$ 3, a weight decay of $81e-$ 8, for 10 epochs, and the mean square error (MSE) as the loss. The IF method was implemented with 100 estimators and $max\_features$ set to one, without bootstrapping. For comparison, the contamination parameter in the IF model was set to match the false positive rate (FPR) obtained by the AE according to the threshold defined in Eq. 1, corresponding to the proportion of benign samples with a reconstruction error above this threshold.
Other anomaly detection methods can be considered depending on the dataset characteristics and attack patterns.
Model Interpretation Stability. In addition to measuring distinguishability, we assess how adversarial samples influence the target model’s decision-making process. Adversarial perturbations may alter the model’s perception of feature importance, leading to inconsistencies in the model’s reasoning. To quantify this effect, we employ SHAP [33], a widely used explainability technique that provides a unified measure of feature importance in ML models. SHAP estimates each feature’s contribution to a prediction by comparing actual predictions with averages across all possible feature combinations.
@@
We introduce two metrics based on SHAP: (1) Importance-Based Anomaly Detection Rate – The percentage of adversarial samples where at least one feature’s SHAP value falls outside the range observed in benign samples, and (2) Average Anomalous SHAP Features per Sample – The average number of features in an adversarial sample with SHAP values outside the benign distribution. Lower values for these metrics indicate higher-quality adversarial samples, as they remain consistent with the model’s internal reasoning despite altering predictions.
@@
Importantly, CSAD is also being extended to SHAP-based anomaly detection; since we assess traditional anomaly detection methods per class, we compute SHAP-based anomalies within the context of each class. This ensures that SHAP value deviations are analyzed relative to the specific class distribution, rather than across the entire dataset. The SHAP values were extracted using SHAP TreeExplainer.
@@ 4 Experimental Setup
4.1 Research Questions
In this study, we address the following research questions regarding the tabular domain:
  • RQ#1 To what extent do query-based attacks reflect the behavior of transferability-based gradient attacks?
  • RQ#2 How the perturbed samples of different attack strategies differ in terms of the attacker’s risk and effort?
  • RQ#3 Is the model’s decision-making process aligned according to the samples’ SHAP values for both a benign samples and perturbed samples?
@@ 4.2 Datasets
    Video Transmission Quality (VideoTQ) dataset. This proprietary commercial dataset was obtained from RADCOM222radcom.com/, a company specializing in service assurance for telecom operators, and it contains records of 54,825 streaming service media broadcasts. This dataset is used to predict a transmission’s resolution (i.e., high or low resolution). The dataset is unbalanced and includes 10,000 records from class 1 (i.e., high resolution). The dataset comprises 23 numerical and categorical features, nine of which are mutable.
@@ 4.3 Dataset Preprocessing
The preprocessing performed on each dataset (illustrated in Fig. 1 (b)) differs according to its characteristics, with the aim of creating highly accurate target models that are based on multiple features’ importance in the model’s decision-making process.
Hate dataset. In the preprocessing performed on the Hate dataset, we followed the preprocessing methodology outlined by Kaggle333kaggle.com/code/binaicrai/fork-of-fork-of-wids-lgbm-gs. We also performed additional preprocessing steps, including removing the ‘user_id’ feature, systematically eliminating outliers, dropping features with a correlation exceeding $90\%$, and excluding instances with more than $80\%$ missing data. In addition, features for which over $75\%$ of the values were the same were removed.
We also performed label encoding on the binary and categorical features and filled in missing values by employing a regression model to predict the values for each feature. Additionally, features that contain the GloVe representations [41] of other features (which include ‘glove’ in their names) were removed, resulting in a distilled set of 311 features. To ensure that only meaningful and important features were used, we trained four models on the dataset (XGBoost [10], GradientBoost [17], LightGBM [27], and random forest [5]) and selected the top-40 important features for each model; the importance of all of these features exceeded 0.0001. The resulting dataset contains 115 features and 4,971 samples.
ICU dataset. In the preprocessing performed on the ICU dataset, we followed the preprocessing methodology outlined by Kaggle444kaggle.com/code/binaicrai/fork-of-fork-of-wids-lgbm-gs. We also performed additional preprocessing, removing outliers, eliminating highly correlated features (with over $90\%$ correlation), and excluding instances with over $80\%$ missing values. In addition, features for which over $75\%$ of the values were the same were removed. Samples for which the ‘bmi’ feature was missing, values were filled in by using the BMI formula based on the values in the ‘height’ and ‘weight’ features. For other features, we filled in missing values by using a regression model to predict the values for each feature. Categorical features were label-encoded, and missing values were filled in by using the median feature value.
@@
VideoTQ dataset. In the preprocessing performed on the VideoTQ dataset, we removed 15 records that had missing values, and eliminated outliers and highly correlated features (with over $90\%$ correlation). Categorical features were label-encoded, and all timestamps were converted into separate date and time formats.
@@
For all datasets, the editable and immutable features were determined and selected by a domain expert. All datasets were split into training and test sets, with $75\%$ of the samples serving as the training set. Oversampling was performed by duplicating the training set samples in class 1 to address the significant class imbalance in the datasets. The oversampled training set was split into a target training set and a surrogate training set (for the attacker’s use in the transferability-based attacks). The test set is also used to create adversarial samples (i.e., the attack set) as illustrated in Fig. 1 (c) and further detailed in Section 4.8.
4.4 Target Models
For our evaluation, we trained four different ML target models: XGBoost [10] (XGB), GradientBoost [17] (GB), LightGBM [27] (LGB), and random forest [5] (RF) (see 1 (d)). The hyperparameters used to train the models were fine-tuned, and therefore varied across different training sets. The XGBoost target models were trained using 90, 70, and 300 estimators; a max depth of 3, 8, and 5; and a learning rate of 1, 0.1, and 1 for the Hate, ICU, and VideoTQ datasets respectively. The GradientBoost target models were trained using 40, 500, and 300 estimators; a max depth of 7, 6, and 5; and a learning rate of 2.5, 0.01, and 1 for the Hate, ICU, and VideoTQ datasets respectively. The LightGBM target models were trained using 300, 200, and 200 estimators; a max depth of 7, 8, and 8; and a learning rate of 1, 0.1, and 0.1 for the Hate, ICU, and VideoTQ datasets respectively. The random forest target models were trained using 100, 500, and 500 estimators and a max depth of 4, 9, and 9 for the Hate, ICU, and VideoTQ datasets respectively.

### v3

@@ 2.2 Evaluation Challenges and Anomaly Detection
Traditional anomaly detection methods developed for image domains often prove inadequate for tabular data. Convolutional autoencoders (CAEs) and generative adversarial networks (GANs), while effective for spatial correlations, struggle to capture the complex multivariate dependencies characteristic of tabular data [40]. Classical statistical approaches like t-tests and the ANOVA (analysis of variance) assume normality and independence, limiting their applicability to tabular data domains where variables might not follow normal distributions and exhibit complex, non-linear relationships[41]. Simple univariate approaches, such as z-score (standard score) thresholding, focus on individual feature deviations but fail to account for multivariate dependencies [42]. Similarly, principal component analysis (PCA)-based methods assume linear relationships, limiting their effectiveness in detecting non-linear dependencies in high-dimensional datasets[42].
@@
More suitable approaches for tabular data include autoencoders (AEs), which learn latent representations to identify deviations from expected patterns [42], and isolation forest (IF) algorithms, which detect anomalies by isolating points in a partitioned feature space. However, while IF excels at detecting structurally shifted or distant outliers, it may exhibit reduced sensitivity to some subtle inconsistencies that may characterize sophisticated adversarial attacks [43].
It should be noted that the anomaly detection framework proposed here serves as a methodological foundation for assessing adversarial sample quality and can be extended to incorporate additional anomaly detection techniques as needed. Alternative approaches such as support vector machines (SVMs) with one-class classification, local outlier factor (LOF) methods, or more recent deep learning-based anomaly detection techniques could similarly be integrated into this evaluation framework. The choice of specific anomaly detection methods can be adapted based on dataset characteristics, computational constraints, and the particular nature of adversarial attacks under investigation, providing flexibility for future research applications.
In this study, IF and AEs were selected as primary anomaly detection baselines, motivated by their complementary strengths and established effectiveness in the tabular domain. IF provides computational efficiency and strong performance on high-dimensional data without requiring distributional assumptions, making it suitable for diverse tabular datasets. Conversely, AEs can model complex feature interactions and distributional shifts that may characterize sophisticated adversarial perturbations. This dual approach enables the comprehensive evaluation of adversarial sample quality across different types of anomalous patterns.
2.3 Research Gap and Contributions.
Table˜1 summarizes related studies that developed adversarial attacks, taking into account the unique constraints of the tabular domain. Despite significant advances in tabular adversarial attacks, prior research shares three critical limitations that this work addresses:
Table 1: Related works that developed adversarial attacks, taking into account the unique constraints of the tabular domain.
Paper
(Attack Type)
Key Contribution Evaluation Metrics Limitation
[6]
(White-box)
Incorporates feature validity constraints and minimizes perturbations on important features to reduce imperceptibility Success rate,
weighted $L_{2}$ norm (scaled by feature importance),
distance to nearest neighbor
Drops non-ordered categorical features; no evaluation of feature coherence
[7]
(Black-box)
Extends the weighted $L_{2}$ norm from [6] to minimize perturbations on both important and unimportant (irrelevant) features; adapts query-based attacks with validity constraints and custom thresholds for unbalanced datasets Success rate,
important feature change count,
validity satisfaction
Limited to validity constraints; no coherence evaluation
[8]
(White-box, Transferability)
Uses NN embeddings to minimize $L_{2}$ distance while preserving feature correlations Success rate,
$L_{0}$ distance
No evaluation of semantic coherence
[9]
(White-box, Transferability)
Extends previous work by incorporating feature importance for perturbation selection, reducing $L_{0}$ distance Success rate,
$L_{0}$ distance
No evaluation of semantic coherence
[31]
Develops a transferability-based approach considering feature distributions across different models Success rate,
transferability improvement
No validity or coherence constraints; focused solely on transferability performance
Evaluation Framework Limitations: Existing approaches lack objective metrics for assessing adversarial sample quality in the tabular domain, relying on subjective expert judgment or conventional metrics that fail to examine sample coherence and consistency. This work develops comprehensive evaluation criteria based on anomaly detection assessment of sample values and feature importance, providing a principled framework to quantify adversarial sample quality and its impact on model decision-making processes.
Anomaly Detection Limitations: Existing anomaly detection methods assess samples against global data distributions, failing to account for the fact that in tabular data, small adversarial changes may not be anomalous relative to their original class but can be anomalous relative to the target class they impersonate after the attack. We introduce CSAD, which assesses anomalies relative to class-specific distributions rather than global thresholds, better capturing the nuanced nature of tabular adversarial perturbations. This approach is further elaborated in Section˜3.3.
Incomplete Comparative Analysis: Recent research lacks systematic comparison of adversarial attack strategies’ performance factors in the tabular domain. This work provides a comprehensive comparison of attack effectiveness and limitations across multiple attack types and datasets, offering valuable insights into their relative strengths and weaknesses.
By establishing these foundations, this research enables more informed attack design and evaluation development in tabular machine learning domains, contributing novel evaluation methodologies and comprehensive empirical insights to the field.
@@
In our evaluation (see Section˜4.4), we examine two decision-based attacks: the boundary attack [25] and the HopSkipJump attack [26], which were adjusted to address tabular constraints and evaluated by  [7]. We implement these attacks by applying additional constraints based on feature distributions (e.g., limiting values according to statistical information) and making adjustments to ensure validity (e.g., [0,1] values for binary features and ordinal values for categorical features) and editability (guided by domain knowledge).
In addition, we perform transferability-based gradient attacks based on the architecture proposed in [8, 9], which is a sophisticated attack designed to craft valid and consistent adversarial samples.
@@
To ensure coherence and maintain feature consistency when crafting adversarial samples, we introduce a novel technique for perturbing dependent features. Specifically, for each dependent feature, we train a regression model on the benign samples, excluding the target feature itself. This model then predicts the value of the target feature based on the values of the other features in the modified sample.
During the adversarial attack process, when perturbations are applied to each dependent feature, the corresponding regression model is queried to predict its updated value, ensuring that the dependent feature remains consistent with the rest of the modified sample. This approach ensures that the perturbations respect consistency between features, thereby maintaining the overall coherence of the sample.
This method addresses scenarios where feature dependencies are known to exist, but the exact computational relationship is either unavailable to the attacker or only partially observable in the dataset. For instance, features such as ‘apache_3j_bodysystem’ and ‘apache_3j_diagnosis’ in the ICU dataset are components of clinical scoring systems (e.g., APACHE - Acute Physiology and Chronic Health Evaluation). These features are derived from multiple physiological and clinical attributes, some of which are not included in the dataset, creating semantic dependencies on other inputs through complex and often unknown relationships. Since the attacker cannot reproduce the precise calculations used in practice, the regression-based approach provides an effective approximation.
When the dependency structure is unknown or latent, an additional preliminary analysis stage is required to identify these relationships before applying the regression-based adjustment. Such identification can be performed using statistical dependency measures, feature clustering, or model-based approaches (e.g., probabilistic graphical models). Without this step, the regression-based perturbation technique cannot be reliably applied.
The framework is inherently modular and allows flexibility in the regression model selection based on dataset characteristics. For datasets with linear feature relationships, linear regression provides computational efficiency with satisfactory predictive accuracy. When complex, non-linear relationships exist, more sophisticated models such as polynomial regression or random forests can be employed to better capture feature interactions, although at increased computational cost. Any suitable dependency modeling technique can be substituted for the regression step, including non-linear or ensemble-based models, to better capture complex interactions and adapt to datasets with complex correlation structures.
This technique focuses on cases where features have complex or unknown dependencies on other variables, while assuming that the features with dependencies are not directly interdependent among themselves. More challenging scenarios arise when multiple features depend on each other simultaneously, forming mutual or circular dependencies. Handling such cases would require an additional methodology to identify the existence of these circular dependencies, determine the specific group of mutually dependent features, and develop solutions that enable coordinated modification of multiple interdependent features; this is an open challenge that remains for future research.
The application of this technique should be selective, based on the characteristics of the dataset. In our evaluation, we identified clear feature dependencies in the ICU dataset that warranted this approach. However, for datasets like the Hateful Users on Twitter and the Video Transmission Quality, our domain analysis revealed no features that are strictly dependent on other unknown or unobserved variables, making the regression-based adjustment unnecessary for these cases.
In our implementation of both query- and transferability-based attack strategies, the attack process includes an additional step in which the dependent features, as identified by domain knowledge, are adjusted according to the predictions from the regression models. This step helps maintain feature consistency, improve realism, and reduce the likelihood of generating detectable anomalies.
@@ 3.3 Evaluating the Quality of Adversarial Attacks
While Feature Space Coherency captures statistical deviations from the benign distribution, Model Interpretation Stability provides a complementary perspective by examining how adversarial samples influence the model’s internal reasoning. This aspect is particularly important in cases where adversarial perturbations are subtle and remain close to the benign data in the input space yet still cause shifts in the model’s attribution of feature importance. Such changes may not be detectable through input-level analysis alone.
@@
In addition to its conceptual advantages, the CSAD approach is computationally efficient in both the training and inference phases. During inference, CSAD maintains constant computational complexity, as each sample is evaluated by a single class-specific model, identical to traditional approaches. The training phase reveals distinct computational advantages that depend both on the computational complexity of the underlying anomaly detection algorithm and on the class distribution in the dataset.
For algorithms with super-linear complexity, i.e., complexity of order $O(n^{\alpha})$ with $\alpha>1$, such as one-class SVM with RBF kernels ($O(n^{2})$ to $O(n^{3})$[44], CSAD can reduce computational overhead compared to traditional approaches that train a single model on the full dataset. In the balanced case, where each of the $k$ classes contains $\approx n/k$ samples, training $k$ models requires $O(k\cdot(n/k)^{\alpha})=O(k^{1-\alpha}\cdot n^{\alpha})$ operations for $\alpha\geq 1$, which is lower than the $O(n^{\alpha})$ cost of the traditional approach for $\alpha>1$. In unbalanced datasets, the computational cost improvement decreases as class imbalance grows, and in the extreme case where one class dominates (i.e., contains most of the samples), the computational cost becomes equivalent to that of the traditional method. A detailed derivation is provided in Appendix B.
For algorithms with linear training complexity, such as IF (O(n)O(n)[45] and AEs (O(n)O(n)[42], partitioning the data into $k$ class-specific subsets does not change the total cost in either balanced or imbalanced scenarios, since the per class training cost scales linearly with the subset size. Training these $k$ models requires $O(n)$ operations in total, matching the computational complexity of the traditional single model approach. CSAD therefore maintains computational parity for linear models while retaining its conceptual advantages for class-specific anomaly detection. Nevertheless, its applicability is inherently constrained by the sample size requirements of the underlying anomaly detection method.
While this work focuses on tabular data, the core principle of CSAD—evaluating anomalies relative to the distribution of benign samples within the predicted class—is general and potentially applicable to other data modalities such as images, text, or graphs. However, extending CSAD to these domains requires the selection and adaptation of anomaly detection methods that align with the specific characteristics and constraints of the data. For example, in tabular data, subtle perturbations may remain close to the original distribution but become anomalous for the predicted class, while in other domains, the notions of anomaly and feature dependencies may differ. Additionally, the ensemble or consensus approach used in CSAD for tabular data—where separate anomaly detection models are trained per class—may not always provide similar benefits in other domains. In some cases, especially with less complex or more homogeneous datasets, a single global anomaly detection model might suffice, as in traditional approaches. Evaluating when and how class-specific modeling improves detection performance in different data modalities remains an open question that must be addressed before expanding CSAD application beyond tabular data.
@@
In the IF algorithm, a sample is considered anomalous if it is classified as an outlier. In the AE model, an anomaly is identified when its reconstruction error exceeds a threshold, defined in Eq.˜1:
$\begin{split}Threshold(x_{c},\>AE_{c})=mean(test_{c,\>r.\,error})\,+\\ 2*std(test_{c,\>r.\,error})\end{split}$ (1)
where $x_{c}$ is a sample classified as class $c$, $AE_{c}$ is the AE trained on benign samples of class $c$, and $test_{c,r.,error}$ represents the reconstruction error on a validation set of class $c$ during the AE’s training. This thresholding approach follows a well-established practice in reconstruction-based anomaly detection using AEs [46, 47, 48]. Such formulations aim to separate benign and anomalous samples based on the statistical distribution of reconstruction errors. Following this approach, we adopt $k=2$, based on empirical tuning to balance sensitivity and false positives across datasets and attack scenarios.
@@
The AE model has a symmetric architecture with a single hidden layer of 64 neurons serving as the embedding layer between the encoder and decoder and ReLU activations. The model was trained using the Adam optimizer with a learning rate of 1e31e-3, a weight decay of 1e81e-8, for 10 epochs, and the mean square error (MSE) loss. This design aligns with common practices in reconstruction-based anomaly detection, where lightweight models have demonstrated good performance with reduced computational demands and lower risk of overfitting, thereby contributing to training stability [49]. The IF method was implemented with 100 estimators and max_featuresmax\_features set to one, without bootstrapping. For comparison, the contamination parameter in the IF model was set to match the false positive rate (FPR) obtained by the AE according to the threshold defined in Eq.˜1, corresponding to the proportion of benign samples with a reconstruction error above this threshold. Other anomaly detection methods can be considered depending on the dataset characteristics and attack patterns.
Model Interpretation Stability. To quantify interpretation stability, we assess how adversarial samples influence the target model’s decision-making process using SHAP [21], a widely used explainability technique that provides a unified measure of feature importance in ML models. SHAP estimates each feature’s contribution to a prediction, computed from the attacked model’s outputs. Specifically, for each class $c$ and feature $f$, we define the normal range as $\min(\text{SHAP}_{f,c}),\max(\text{SHAP}_{f,c})]$ based on the SHAP values observed in benign training samples assigned to class $c$. An adversarial sample predicted as class $c$ is considered to exhibit an anomalous SHAP value for feature $f$ if its value falls outside this class-specific range.
@@
Based on this definition, we introduce two SHAP-based importance metrics: (1) Importance-Based Anomaly Detection Rate – the percentage of adversarial samples where at least one feature’s SHAP value falls outside the benign range (i.e., is considered anomalous); and (2) Average Anomalous SHAP Features per Sample – the average number of features per adversarial sample whose SHAP values fall outside the benign range. Lower values for these metrics indicate higher-quality adversarial samples, as they remain consistent with the model’s internal reasoning despite altering predictions.
@@
Importantly, we have extened CSAD to to SHAP-based anomaly detection; we compute SHAP-based anomalies within the context of each class, ensuring that SHAP value deviations are analyzed relative to the specific class distribution, rather than across the entire dataset. The SHAP values were extracted using SHAP TreeExplainer.
@@ 4.1 Datasets
    Video Transmission Quality (VideoTQ) dataset. This proprietary commercial dataset was obtained from RADCOM,222radcom.com/ a company specializing in service assurance for telecom operators, and it contains records of 54,825 streaming service media broadcasts. This dataset is used to predict a transmission’s resolution (i.e., high or low resolution). The dataset is unbalanced and includes 10,000 records from class 1 (i.e., high resolution). The dataset comprises 23 numerical and categorical features, nine of which are mutable.
@@ 4.2 Dataset Preprocessing
The basic preprocessing pipeline was designed following the established methodology of Mathov et al. [8], which defines standard and widely accepted preprocessing practices for tabular data. This approach was deliberately chosen to ensure fair comparison between attack methods, given our focus on evaluating and comparing adversarial attack behaviors and the coherence of the adversarial samples they generate. By maintaining a consistent preprocessing pipeline across all datasets, we ensure that any variations in results are attributable to the attack behavior itself rather than inconsistencies in data preparation. This standardized approach also facilitates direct comparability with previous works that used similar preprocessing on the same datasets. Additional dataset-specific adaptations were applied where appropriate, to accommodate unique characteristics while preserving the overall consistency of the evaluation framework.
In accordance with the implementation of a standard preprocessing pipeline across all datasets, we systematically eliminated outliers, dropped a group of features with a correlation exceeding $90\%$ and randomly included just one of them, and excluded instances with more than $80\%$ missing data. In addition, unique features, as well as features for which over $75\%$ of the values were identical, were removed. Categorical features were label-encoded, following the methodology of Mathov et al. [8] and Grolman et al. [9]. While one-hot encoding is the standard approach for nominal features, as it preserves equal semantic distance between categories, these works opted for label encoding due to constraints imposed by adversarial attack requirements. Specifically, applying perturbations to one-hot encoded features can lead to invalid states, where multiple elements of the same categorical feature may simultaneously flip their value from zero to one, violating the inherent constraint of mutual exclusivity and resulting in unrealistic adversarial samples. Label encoding maintains the single feature perturbation paradigm, ensuring that adversarial perturbations generate semantically valid examples [8, 9].
The following additional preprocessing steps were applied based on the unique characteristics of each dataset:
For the Hate dataset, we followed the methodological procedure of Grolman et al. [9], which included removing the ‘user_id’ feature and excluding features that contain GloVe representations [50] of other features (which include ‘glove’ in their names), resulting in a distilled set of 311 features. Following the same methodology, we trained four models on the dataset: XGBoost [51], gradient boosting [52], LightGBM [53], and random forest [54], and selected the top-40 most important features for each model based on the Pearson correlation coefficient, ensuring that only meaningful and relevant features were retained. All of the selected features had importance scores exceeding 0.0001, whereas the remaining features exhibited near-zero importance scores. The final processed dataset consists of 115 features and 4,971 samples.
For the ICU dataset, we followed the preprocessing methodology outlined in a publicly available Kaggle implementation.333kaggle.com/code/binaicrai/fork-of-fork-of-wids-lgbm-gs This included both the basic preprocessing steps described above and additional steps proposed in the Kaggle solution. Specifically, the training and test sets provided by Kaggle were combined to increase the amount of available data, as the original training set contained a substantial proportion of missing values, while the test set was comparatively more complete. Missing values were handled through a multi-step imputation process. Samples for which the ‘bmi’ feature was missing were filled in using the BMI formula based on the values in the ‘height’ and ‘weight’ features. Missing values in the following features were imputed using regression models to predict values for each feature: ‘age,’ ‘height,’ ‘weight,’ ‘apache_4a_hospital_death_prob,’ and ‘apache_4a_icu_death_prob.’ Following the imputation process described in Mathov et al. (2022), the remaining numeric features were imputed using median values, and categorical features were imputed using modal values to avoid filling in invalid or non-integer values and ensure consistency with the original feature distribution.
We acknowledge the limitations of using the Kaggle-based preprocessing, specifically the potential introduction of bias due to merging the training and test sets. This approach may alter the data distribution and affect class balance; however, given that the original training set was already highly unbalanced and contained substantial missing data, we considered this trade-off acceptable. Furthermore, the selected preprocessing pipeline demonstrated strong classification performance in the Kaggle setting on the same task as in this paper, supporting its suitability for our objective.
@@
For the VideoTQ dataset, we removed 15 records that had missing values and converted all timestamp features into separate date and time formats.
@@
For all datasets, the editable and immutable features were determined and selected by a domain expert. All datasets were split into training and test sets, with $75\%$ of the samples serving as the training set. Oversampling was performed by duplicating the training set samples in class 1 to address the significant class imbalance in the datasets. The oversampled training set was split into a target training set and a surrogate training set (for the attacker’s use in the transferability-based attacks). The test set is also used to create adversarial samples (i.e., the attack set) as illustrated in Fig.˜1 and further detailed in Section˜4.7.
4.3 Classification Models Used for Evaluation
For our evaluation, we trained four different ML classification models commonly used in real-world tabular data: XGBoost [51] (XGB), gradient boosting [52] (GB), LightGBM [53] (LGB), and random forest [54] (RF) (see Fig.˜1). The hyperparameters for each model were manually tuned to fulfill two key objectives: (i) to achieve high classification performance on each dataset, and (ii) to promote model robustness by encouraging the model to rely on a diverse set of features instead of overfitting to a small number of highly correlated features. The XGBoost target models were trained using 90, 70, and 300 estimators; a max depth of 3, 8, and 5; and a learning rate of 1, 0.1, and 1 for the Hate, ICU, and VideoTQ datasets respectively. The gradient boosting target models were trained using 40, 500, and 300 estimators; a max depth of 7, 6, and 5; and a learning rate of 2.5, 0.01, and 1 for the Hate, ICU, and VideoTQ datasets respectively. The LightGBM target models were trained using 300, 200, and 200 estimators; a max depth of 7, 8, and 8; and a learning rate of 1, 0.1, and 0.1 for the Hate, ICU, and VideoTQ datasets respectively. The random forest target models were trained using 100, 500, and 500 estimators and a max depth of 4, 9, and 9 for the Hate, ICU, and VideoTQ datasets respectively.